the width of recangle is 5 feet longer than one forth of its lenght. the perimeter is 80 feet. find its dimensions
Answer by rfer(16322)

2x+2(0.25x+5)=80
2x+0.5x+10=80
2.5x=80-10
2.5x=70
x=70/2.5
x=28ft
0.25x+5=12ft